OpenStudy (omgthatgirl):

Help Ty has 230 pages of text to read in the next three days. If he averages 25 pages every 1 1/4 hours, how many hours is he going to be spending reading this text? A) 9.2 hours B) 9.5 hours C) 10.25 hours D) 11.5 hours Calculate the average rate of change of f(x) = x2 - 1x - 4 for 2 ≤ x ≤ 6. A) -47/6 B) -6/47 C) 47/6 D) 6/47

OpenStudy (misty1212):

first one we can do a couple different ways maybe easiest to calculate the rate in pages per hours, then divide 230 by that number

OpenStudy (misty1212):

\[25\div 1.25=20\] so reads 20 pages per hour

OpenStudy (misty1212):

then \[230\div 20=11.5\] so 11 and a half hours

OpenStudy (misty1212):

for the second one \[f(x) = x^2 - 1x - 4\] we want the average rate of change from 2 to 6 in other words we have to compute \[\frac{f(6)-f(2)}{6-2}\]

OpenStudy (omgthatgirl):

Of course. I got It now. It is just that i forget how to do these. Thank you so much

OpenStudy (misty1212):

\[f(6)=6^2-6-4=36-6-4=26\] \[f(4)=4^2-4-4=16-8=8\]

OpenStudy (misty1212):

then compute \[\frac{26-8}{6-2}\] and that is all

OpenStudy (omgthatgirl):

Thanks